what is the average velocity of a nitrogen molecule at 28 degrees C and What is the average kinetic energy of a nitrogen molecule at 28 degrees C. what are the units?
Answer #1

Average Kinetic Energy of molecule is given by KE = 3/2kT

k is the Boltzmann constant = 1.38 x 10-23 J/K

T = 28 C = 273+28 = 301 K

Avg KE = 1.5 x 1.38 x 10-23 J/K x 301 K = 6.23 x 10-21 J

KE = 1/2xmxV2

m = mass of nitrogent = 28 amu = 4.65 x 10-26 Kg

V = sqrt(2 x KE/m) = sqrt(2x6.23 x 10-21 J/4.65 x 10-26 Kg) = 517.64 m/s